OpenStudy (anonymous):

A calorie is a unit of heat or energy and it is equal to about 4.2J where 1J=1ksm^2s^-2.Suppose we employ a system of units in which the unit of mass equals skg,the unit of lenght =rm,the unit of time=gs.show that a calorie has a magnitude 4.2 s^-1r^-2g^2.

OpenStudy (anonymous):

Given that, 1 calorie = 4.2 (1 kg) (1 m2) (1 s–2) New unit of mass = α kg Hence, in terms of the new unit, 1 kg = In terms of the new unit of length, And, in terms of the new unit of time, ∴1 calorie = 4.2 (1 α–1) (1 β–2) (1 γ2) = 4.2 α–1 β–2 γ2